OpenStudy (anonymous):

Tell whether or not the following statement are ture for all the numbers. If the statement is not true, give a counterexample. if a>b, and c is positive, then a/c>b/c. help!!!!

OpenStudy (anonymous):

okk its true.. can u xplain me?

OpenStudy (anonymous):

well if a is positive and b is negative it would make the statement true am i right?

OpenStudy (anonymous):

it will in any case even if a and b both positive or if a and b both negatv or one positive and other negatv

OpenStudy (anonymous):

ok thnx!! alot

OpenStudy (anonymous):

Given \[a>b\] assume \[\frac{a}{c}<\frac{b}{c}\]and arrive at a contradiction

OpenStudy (anonymous):

wait clarify "all the numbers"

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

@ebbflo how can it become less than as c is positive number?

OpenStudy (anonymous):

@vedic exactly, that's the contradiction i.e. given \[a>b\] assume \[\frac{a}{c}<\frac{b}{c}\] multiply both sides by \[c\] to get \[a<b\] which is a contradiction thus \[\frac{a}{c}>\frac{b}{c}\]